OpenStudy (anonymous):

How do I factor 8y^2 - 10xy / 4y? Please help me answer this question..

jimthompson5910 (jim_thompson5910):

\[\Large \frac{8y^2 - 10xy}{4y}\] \[\Large \frac{2y(4y - 5x)}{4y}\] \[\Large \frac{2y(4y - 5x)}{2*2y}\] \[\Large \frac{\cancel{2y}(4y - 5x)}{2*\cancel{2y}}\] \[\Large \frac{4y - 5x}{2}\] So \[\Large \frac{8y^2 - 10xy}{4y} = \frac{4y - 5x}{2}\]